OpenStudy (anonymous):

I need help with factoring polynomials by grouping. Here's an example. 3n^3 – 6n^2 + 4n – 8

OpenStudy (mathstudent55):

Find the GCF of the first two terms and factor it out. Find the GCF of the last two terms and factor it out.

OpenStudy (anonymous):

When finding the GCF of 3n^3 – 12n^2, my book says the answer is 3n^2(n – 4). I understand the 3n^2, but where does the (n - 4) come from?

OpenStudy (mathstudent55):

Factoring is the opposite of multiplying out. Look at this simple example. Factor x^2 - x. The GCF is x, so an x comes out. x( ) Now you need to put inside the parentheses the quantities that when multiplied by x will give you back x^2 - x. x( ) = x^2 - x What times the x outside will give you x^2? An x, so an x goes inside the parentheses: x(x ) = x^2 - x Now what times the x outside will give you -x? -1, so the -1 goes inside the parentheses. x(x - 1) = x^2 - x Check: multiply x(x - 1) by distributing the x: x^2 - x, so factoring is correct.

OpenStudy (mathstudent55):

Now let's look at your problem. The first two terms are 3n^3 - 6n^2, and you understand the GCF is 3n^2. The factorization is going to end up as: 3n^2( ) What do you need to multiply 3n^2 by to get 3n^3? Answer: n 3n^2(n ) What do you need to multiply 3n^2 by to get -6n^2? Answer: -2 3n^2(n - 2)

OpenStudy (mathstudent55):

Your question: When finding the GCF of 3n^3 – 12n^2, my book says the answer is 3n^2(n – 4). I understand the 3n^2, but where does the (n - 4) come from? See above answer. It's similar. 3n^3 - 12n = 3n^2( ) 3n^2 times what equals 3n^3? Answer: n, so 3n^3 - 12n = 3n^2(n ) 3n^2 times what equals -12n? Answer: -4, so 3n^3 - 12n = 3n^2(n - 4)

OpenStudy (anonymous):

Thank you. My mathbook showed the answers to the example problem but for some reason didn't really show how to get the answers. I understand it now.

OpenStudy (mathstudent55):

Here's the rest of the factorization: 3n^3 – 6n^2 + 4n – 8 3n^2(n - 2) +4(n - 2) (n - 2)(3n^2 + 4)